OpenStudy (anonymous):

How can I calculate the enthalpy combustion of hydrocarbon? Please help me out

OpenStudy (anonymous):

@ganeshie8 ?

OpenStudy (tatianagomezb):

the enthalpy of combustion for certain substances can be thought of as the amount of energy required to "burn" 1 mole of said substance. When you say Hydrocarbon which one do you mean? You can calculate the enthalpy of any reaction as follows: \[\Delta H reaction= \sum n \Delta H _{f} products - \sum n \Delta H _{f} reactants \] Where \[\Delta H \] is the enthalpy of formation in standard conditions and should be available in a table given to you by professors, and n is the numbers of moles. Les think about this reaction \[CH _{4} (g) + 2 O _{2} (g) \rightarrow CO _{2} (g) + 2 H _{2}O (l) \] the enthalpy of combustion of methane gas with oxygen can be written as \[\Delta H rxn= [(2 mol \times \Delta H _{f} H _{2}O (l)) + (1 mol \times \Delta H _{f} CO _{2})]-\left[ 1 mol \times \Delta H _{f} CH _{4} \right]\] Elements such as oxygen gas have an enthalpy of formation of 0, that's why it's not included. All you have to do is find the enthalpies of formation for each substance, and keep in mind that depending on the reaction, heat is absorbed or released to the system.